# Learning Partial Differential Equations with Deep Parallel Neural Operators

### v1

@@ 1 Introduction
Fourier neural operators have proposed a novel solution that performs well in handling partial differential equation (PDE) problems. It transforms the data into the Fourier domain for learning [9]. However, Fourier neural operators rely on the iterative update strategy of the operator, which can result in partial information loss during iterative calls. The main reason for this error accumulation is the frequent conversion between the spatial and spectral domains within different operational blocks, as well as negative multiplications performed inside the Fourier domain. In this process, some spatial information is lost due to the forward and backward transformations of spectral convolution. Therefore, the cumulative loss of spatial information in processing the same information after each iteration reduces its efficiency [10, 11].
@@ 2.2 Fourier transform and Fourier neural operator
In the Fourier domain, differential operations become simpler because differentiation is equivalent to multiplication in the Fourier domain. This makes Fourier transform a powerful tool for solving partial differential equations (PDEs), especially in cases involving frequency components. Through Fourier transform, complex differential equations can be transformed into algebraic equations that are easier to handle, thereby simplifying the solving process. Fourier transform has played an important role in the development of deep learning. In theory, Fourier transform appears in the proof of the general approximation theorem, which states that multi-layer neural networks can approximate any continuous function [19]. In practice, they are often used to accelerate convolutional neural networks [20]. Neural network architectures associated with Fourier transforms or sinusoidal activation functions are proposed and studied [21, 22, 23].
@@ 4.1.1 Navier-stokes equation
Where $\nu$ is the fixed viscosity, $u$ denotes the velocity, $p$ is the pressure, and $f$ is the external force function. The dataset is provided by the baseline model, and we used data with a grid resolution of $64\times 64$ for both training and testing.
@@ 4.1.2 Plastic problem
The dataset is from the baseline model geo-FNO. Contains 980 training data and 80 test data. All data points are contained on a 101 × 31 structured grid with 20 time steps.
4.1.3 Other datasets
Other datasets include the Darcy flow problem, the elastic problem, the airfoil problem for the euler equations, and the pipe problem for the Navier-Stokes equations. Darcy flow problem, We consider the steady state of the two-dimensional Darcy flow equation in the unit rectangle. This is a second order linear elliptic partial differential equation with Dirichlet boundary conditions experiments are performed using a fixed resolution of 85 * 85. The elastic problem is a solid domain problem where the input is mesh information and the target output is the corresponding stress field. The airfoil problem with Euler equations describes the flow characteristics of compressible air around a wing. The inputs are mesh point and Mach number information and the output is the Mach number. The pipe problem using the Navier-Stokes equations describes the flow of an incompressible fluid through a pipe. Inputs include grid point locations and horizontal velocities, and output data include velocity vectors and pressures. The above dataset is provided by the baseline model and consists of one solid-related problem and three fluid-related problems.
4.1.4 Baselines and Settings
In order to better demonstrate the performance of our proposed model, we compare DPNO with seven currently recognized state-of-the-art models, including U-Net, FNO, WMT, U-FNO, UNO, F-FNO, and LSM. Among them, LSM is currently the best overall performer on the six datasets. We used the same experimental setup for all models. The ADAM optimizer was used for all methods, with an initial learning rate of $10^{-3}$ for L2 loss and 501 training sessions. The batch size was set to 20.
@@ 4.2 Results and Discussion
Table 2: Comparison of L2 errors for the 7 benchmark models on the 6 benchmark datasets. The best performing models are in bold, and the next best models are underlined.
@@
The experimental results are visualised in Figures 2-4:
Figure 2: Plastic Problem: The top represents the true value, the middle represents the DPNO prediction, and the bottom line represents the FNO prediction. The five columns represent time variations. Colours represent displacement criteria.
Figure 3: Darcy Flow Problem(left) and Elastic Problem(right): true values on the left, in the middle are the predicted values and errors for DPNO (top) and FNO (bottom)
Figure 4: Airfoil(left) and Pipe Problem(right): top is the true value, middle and bottom are the predicted value and error for DPNO (left) and FNO (right) respectively.
@@ 4.3 Ablation experiment
In order to verify the effectiveness of the parallel strategy in the model, we design a set of ablation experiments to change the parallel block into serial block and observe the performance changes. Specifically, we merge the parallel block horizontally into serial block as shown in Fig.5.
Figure 5: Parallel operator block (a) and serial operator block (b).
@@
Table 3: Comparison of L2 errors for DPNO/P and DPNO on the 6 benchmark datasets. The improvement percentage is included in the DPNO row.
@@ 5 Conclusions
Neural operators are good learners when it comes to functional spaces. In this paper, we propose a deeply parallel Fourier neural operator to address the problem of error accumulation during FNO iterations. Firstly, the data are projected into various potential spaces and learnt in an end-to-end manner through multiple parallel operator blocks. Excellent performance is achieved on six benchmark datasets for partial differential equations. An average improvement of 39% is achieved compared to the baseline model FNO.
Although the field of deep learning currently has good progress in learning partial differential equations. However, there are still shortcomings in long time prediction. The time dependence cannot be captured better. For dynamically changing systems, deep learning models may need to be constantly updated and retrained to adapt to new conditions.

### v2

@@ 1 Introduction
Fourier neural operators have proposed a novel solution that performs well in handling partial differential equation problems. It transforms the data into the Fourier domain for learning [9]. However, Fourier neural operators rely on the iterative update strategy of the operator, which can result in partial information loss during iterative calls. The main reason for this error accumulation is the frequent conversion between the spatial and spectral domains within different operational blocks, as well as negative multiplications performed inside the Fourier domain. In this process, some spatial information is lost due to the forward and backward transformations of spectral convolution. Therefore, the cumulative loss of spatial information in processing the same information after each iteration reduces its efficiency [10, 11].
@@ 2.2 Fourier transform and Fourier neural operator
In the Fourier domain, differential operations become simpler because differentiation is equivalent to multiplication in the Fourier domain. This makes Fourier transform a powerful tool for solving partial differential equations, especially in cases involving frequency components. Through Fourier transform, complex differential equations can be transformed into algebraic equations that are easier to handle, thereby simplifying the solving process. Fourier transform has played an important role in the development of deep learning. In theory, Fourier transform appears in the proof of the general approximation theorem, which states that multi-layer neural networks can approximate any continuous function [19]. In practice, they are often used to accelerate convolutional neural networks [20]. Neural network architectures associated with Fourier transforms or sinusoidal activation functions are proposed and studied [21, 22, 23].
@@ 4 Experiment and Results
4.1 Baselines and Settings
In order to better demonstrate the performance of our proposed model, we compare DPNO with seven currently recognized state-of-the-art models, including U-Net, FNO, WMT, U-FNO, UNO, F-FNO, and LSM. Among them, LSM is currently the best overall performer on the six datasets. We used the same experimental setup for all models. The detailed experimental setup is shown in Table 2:
@@ 4.1 Baselines and Settings
Table 2: Settings.
epoch batch_size learning_rate optimizer
500 20 0.001 Adam(1e-4)
To ensure that all models converge, we set the number of training rounds to 500. Training batch size of 20. Each parameter of the model will be updated using the Adam optimiser with a learning rate of 0.001 and a weight decay of 0.0001.
4.1.1 Evaluation Metrics
Based on the problem setting, the whole task is to solve for one or more sets of outputs from one set of inputs. The numerical values themselves are small, so in order to describe the model performance more objectively. We use MSE as an evaluation metric.
MSE is a commonly used metric for evaluating the performance of predictive models and is used in various types of machine learning tasks. It reflects the magnitude of the model’s error by measuring the average squared difference between the model’s predicted and actual values.The smaller the value of MSE, the closer the model’s predictions are to the true values.
Given a dataset containing $n$ samples, where each sample has an actual value $y_{i}$ and a corresponding predicted value $\hat{y}_{i}$, the MSE is given by:
$MSE$ $=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (13)
4.2 Datasets and Results
This section will specifically describe the six datasets and show the experimental results of the proposed model and the seven baselines in this paper.
@@ 4.2.1 Navier-stokes Equation
Where $\nu$ is the fixed viscosity, $u$ denotes the velocity, $p$ is the pressure, and $f$ is the external force function. The dataset is provided by the baseline model, and we used data with a grid resolution of $64\times 64$ for both training and testing. The results are shown in Table 3. Detailed visualisation results are in Appendix A.
Table 3: Comparison of Models for Navier-stokes Equation
Model Navier-stokes Equation
U-Net(2015) 0.1982
geo-FNO(2023) 0.1556
WMT(2021) 0.1541
U-NO(2022) 0.1731
U-FNO(2021) 0.2231
F-FNO(2023) 0.2322
LSM(2023) 0.1535
DPNO 0.1418(7.6%)
@@ 4.2.2 Plastic Problem
The dataset is from the baseline model geo-FNO. Contains 980 training data and 80 test data. All data points are contained on a 101 × 31 structured grid with 20 time steps [28]. The results are shown in Table 4 and Fig 2.
(a) Ground True: Each column corresponds to a time point
(b) DPNO Predictions
(c) FNO Predictions
Figure 2: Plastic Problem
Table 4: Comparison of Models for Plasticity
Model Plasticity
U-Net(2015) 0.0051
geo-FNO(2023) 0.0074
WMT(2021) 0.0076
U-NO(2022) 0.0034
U-FNO(2021) 0.0039
F-FNO(2023) 0.0047
LSM(2023) 0.0025
DPNO 0.0019(24%)
4.2.3 Darcy Flow
Henry Darcy’s law describes the flow characteristics of fluids through porous media. This law finds extensive applications across numerous natural science fields, including Fourier’s law in heat conduction, Ohm’s law in electric current, and Fick’s law in diffusion theory. The fundamental form of Darcy’s law is:
$Q$ $=-KA\frac{dh}{dl}$ (16)
Where $Q$ refers to the volume of fluid passing through the medium per unit time, $K$ is the permeability coefficient, $A$ is the flow cross-sectional area, and $\frac{dh}{dl}$ is the pressure gradient. We consider the steady state of the two-dimensional Darcy flow equation in a unit rectangle. It is a second-order, linear, elliptic partial differential equation with Dirichlet boundary conditions. We used the dataset provided in the baseline paper, with a grid size of 421 * 421. Other resolutions were extracted from the highest resolution, and our experiments used a fixed resolution of 85 * 85. The results are shown in Table 5 and Fig 3. To better demonstrate the generalization capability of the model, comparison experiments at different resolutions were added, as shown in Table 6.
Table 5: Comparison of Models for Darcy
Model Darcy
U-Net(2015) 0.0080
geo-FNO(2023) 0.0108
WMT(2021) 0.0082
U-NO(2022) 0.0183
U-FNO(2021) 0.0183
F-FNO(2023) 0.0077
LSM(2023) 0.0065
DPNO 0.0062(4.6%)
Table 6: Comparison of Models for Darcy at Different Resolutions
Model 141*141 211*211 421*421
FNO 0.0109 0.0109 0.0098
UNO 0.0060 0.0058 0.0057
DPNO 0.0057 0.0056 0.0052
(a) DPNO, FNO, and LSM predictions.
(b) Differences in absolute values between the predictions of DPNO, FNO, LSM, and the ground truth.
Figure 3: Darcy Flow.
4.2.4 Elastic Problem
The governing equations related to solid mechanics are described as follows:
$\rho_{s}\frac{\partial^{2}u}{\partial t^{2}}+\nabla\cdot\sigma$ $=0$ (17)
where $\rho_{s}$ is the mass density, $u$ is the displacement vector, and $\sigma$ is the stress tensor. To complete the system, constitutive models are needed to establish the relationship between the strain tensor $\epsilon$ and the stress tensor $\sigma$.
Consider a unit plane $\Omega=[0,1]\times[0,1]$ with an arbitrary shaped void. The bottom of the plane is fixed, and a traction force $t=[0,100]$ is applied at the top. The material is the incompressible Rivlin-Saunders material. The corresponding constitutive model for the material is given by:
$\sigma=\frac{\partial w(\epsilon)}{\partial\epsilon}$ (18)
$w(\epsilon)=C_{1}(I_{1}-3)+C_{2}(I_{2}-3)$ (19)
where $I_{1}$ and $I_{2}$ are scalar invariants of the right Cauchy Green stretch tensor $C=2\epsilon+I$.$C_{1}$ is a tunable energy density parameter. In the experiments, we adopted 1000 training points and 200 testing points [28]. The results are shown in Table 7 and Fig 4.
Table 7: Comparison of Models for Elastic-G
Model Elastic-G
U-Net(2015) 0.0531
geo-FNO(2023) 0.0508
WMT(2021) 0.0520
U-NO(2022) 0.0469
U-FNO(2021) 0.0480
F-FNO(2023) 0.0475
LSM(2023) 0.0408
DPNO 0.0375(8%)
(a) DPNO, FNO, and LSM predictions.
(b) Differences in absolute values between the predictions of DPNO, FNO, LSM, and the ground truth.
Figure 4: Elastic Problem.
4.2.5 Airfoil problem with Euler’s equation
The Airfoil Problem can be analyzed using Euler’s equations, which describe the motion of an inviscid (non-viscous) fluid. These equations are a simplified form of the Navier-Stokes equations and are applicable to the flow around airfoils at high speeds where viscous effects are negligible. Euler’s equations can be described as follows:
$\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\mathbf{u})=0$ (20)
$\frac{\partial(\rho\mathbf{u})}{\partial t}+\nabla\cdot(\rho\mathbf{u}\otimes\mathbf{u})+\nabla p=0$ (21)
$\frac{\partial E}{\partial t}+\nabla\cdot(E\mathbf{u})+p\nabla\cdot\mathbf{u}=0$ (22)
where $\rho$ is the mass per unit volume, reflecting the concentration of the fluid; $\mathbf{u}$ is the distance travelled by a fluid particle per unit time, a vector describing the change in fluid momentum and flow characteristics; $p$ is the force acting per unit area; and $E$ is the total energy of the system. In this case, the far-field boundary conditions are $\rho_{1}=1$, $p_{1}=1.0$, $M_{1}=0.8$, and $\text{AoA}=0$, where $M_{1}$ is the Mach number and AoA is the angle of attack. Additionally, a no-penetration condition is imposed at the airfoil [28]. The experiment uses 1000 training grid points and 200 test grid points. The results are shown in Table 8 and Fig 5.
(a) DPNO, FNO, and LSM predictions.
(b) Differences in absolute values between the predictions of DPNO, FNO, LSM, and the ground truth.
Figure 5: Airfoil problem.
Table 8: Comparison of Models for Airfoil Problem
Model Airfoil
U-Net(2015) 0.0079
geo-FNO(2023) 0.0138
WMT(2021) 0.0075
U-NO(2022) 0.0078
U-FNO(2021) 0.0269
F-FNO(2023) 0.0078
LSM(2023) 0.0059
DPNO 0.0054(8.5%)
4.2.6 Pipe problem with Navier-Stokes equation
The Pipe Problem in fluid dynamics typically refers to the study of fluid flow through a pipe, governed by the Navier-Stokes equations. n the context of the pipe problem, the Navier-Stokes equations can be simplified under the assumption of steady, incompressible, and laminar flow. The simplified forms are as follows:
$\nabla\cdot\mathbf{v}=0$ (23)
$\frac{\partial\mathbf{v}}{\partial t}+(\mathbf{v}\cdot\nabla)\mathbf{v}=-\nabla p+\nu\nabla^{2}\mathbf{v}$ (24)
where $\mathbf{v}$ is the velocity vector, $p$ is the pressure, and $\nu=0.005$ is the viscosity. In this case, at the inlet, a parabolic velocity profile is established with a maximum velocity of $\mathbf{v}=[1;0]$. The outlet is treated with a free boundary condition, while the pipe surface is subject to a no-slip boundary condition. The pipe has a length of 10 and a diameter of 1. The centerline of the pipe is defined using four piecewise cubic polynomials, which are derived from the vertical positions and slopes at five evenly spaced control nodes. The vertical positions at these nodes follow a uniform distribution $d\sim U[-2;2]$, and the slopes at these nodes are also distributed uniformly, $d\sim U[-1;1]$ [28]. The experiment uses 1000 training grid points and 200 test grid points. The results are shown in Table 9 and Fig 6.
Table 9: Comparison of Models for Pipe Problem
Model Pipe
U-Net(2015) 0.0065
geo-FNO(2023) 0.0067
WMT(2021) 0.0077
U-NO(2022) 0.0100
U-FNO(2021) 0.0056
F-FNO(2023) 0.0070
LSM(2023) 0.0050
DPNO 0.0052(-3.8%)
(a) DPNO, FNO, and LSM predictions.
(b) Differences in absolute values between the predictions of DPNO, FNO, LSM, and the ground truth.
Figure 6: Pipe Problem.
@@ 4.3 Results and Discussion
Table 10: Comparison of MSE for the 7 benchmark models on the 6 benchmark datasets. The best performing models are in bold, and the next best models are underlined.
@@ 4.4 Ablation Experiment
In order to verify the effectiveness of the parallel strategy in the model, we design a set of ablation experiments to change the parallel block into serial block and observe the performance changes. Specifically, we merge the parallel block horizontally into serial block as shown in Fig 7.
(a) Parallel Operator Block.
(b) Serial Operator Block.
Figure 7: Comparison of Operator Block Architectures.
@@
Table 11: Comparison of MSE for DPNO/P and DPNO on the 6 benchmark datasets. The improvement percentage is included in the DPNO row.
@@ 5 Conclusions
Neural operators are good learners when it comes to functional spaces. In this paper, we propose a deeply parallel Fourier neural operator to address the problem of error accumulation during FNO iterations. The data are projected into various potential spaces and learnt in an end-to-end manner through multiple parallel operator blocks. Excellent performance is achieved on six benchmark datasets for partial differential equations. An average improvement of 39% is achieved compared to the baseline model FNO.
Although the field of deep learning currently has good progress in learning partial differential equations. However, there are still shortcomings in long time prediction. The time dependence cannot be captured better. For dynamically changing systems, deep learning models may need to be constantly updated and retrained to adapt to new conditions. Therefore, future work tends to enhance the neural operator model by integrating the physical information model. By integrating physical information, the data dependency of the neural operator model is reduced.
Appendix A Navier-stokes Experiment
The following is a visualisation of the experimental results for different models on the N-S equation, as shown in Figs 8.
Where the first row is the model prediction, the second row is the absolute value error. The Navier-stokes equation used in this paper is a task that includes a time term. Each column in the image represents a state at a different point in time.
(a) Experimental results of DPNO on Navier-stokes Equation.
(b) Experimental results of FNO on Navier-stokes Equation
(c) Experimental results of LSM on Navier-stokes Equation.
Figure 8: Experimental results on Navier-stokes Equation.